OpenStudy (anonymous):

sin[cos^-1(2/3)] Find EXACT please

geerky42 (geerky42):

\[\Huge \dfrac{\sqrt{5}}{3}\]

OpenStudy (anonymous):

sorry, i had the answer, would you mind showing how you came to that solution?

geerky42 (geerky42):

Let x be an unknown angle and set it equal to cos^-1(2/3). 2/3 = cos(x) Now draw triangle. Then use Pythagorean theorem to find the "?," then you should know what to do next. Is this clear?

OpenStudy (anonymous):

right on, thanks a lot mate =)

geerky42 (geerky42):

Glad I helped.